The perimeter of a rectangular swimming pool is 154 m. Its length is 2 m more than twice its breadth. What are the length and the breadth of the pool? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given that,
Perimeter of rectangular swimming pool = 154 m
Let the breadth of rectangle be = x According to the question,
Length of the rectangle = 2x + 2 We know that,
Perimeter = 2(length + breadth)
⇒2(2x+2+x)=154m
⇒ 2(3x + 2) = 154
⇒3x+2=154/2
⇒3x=77–2
⇒3x=75
⇒x=75/3
⇒x=25m
Therefore, Breadth = x = 25 cm Length=2x+2
= (2 × 25) + 2
=50+2
=52m

Thanks